When a single slit is replaced bydouble slit arrangement, the intensity of central maxima is increased by?
1
Expert's answer
2020-06-08T10:29:41-0400

We know that the intensity due to single silt

"I_1=I_o(\frac{\sin\phi}{\phi})^2"

Intensity due to double slit

"I_2=I_o \cos^2(\frac{\phi}{2})"

From first and second,

"\frac{I_1}{I_2}=\frac{4\sin^2(\phi/2)}{\phi}"